OpenStudy (anonymous):

I'm study the recursive formula for an arithmetic sequence in algebra 2. Is xn = a + d(n-1) the same as an+1 = an + d ?? ( the an are actually a subscript n)

OpenStudy (anonymous):

yes if i am reading it correctly

OpenStudy (anonymous):

I'm kinda confused about the n subscript thing

OpenStudy (anonymous):

actually i changed my mind the first one has an x and the second one doesn't

OpenStudy (anonymous):

What does the whole subscript thing have to do with this

OpenStudy (anonymous):

X and a r the same

OpenStudy (anonymous):

i am not sure what \[x_n=a+d(n-1)\] means

OpenStudy (anonymous):

Just used different variables for the two equations

OpenStudy (anonymous):

\[a_{n+1}=a_n+d\] is easy to understand it says the next term is \(d\) more than the previous one

OpenStudy (anonymous):

By what's with the n+1

OpenStudy (anonymous):

That's what is confusing me

OpenStudy (anonymous):

lets do an example suppose \(a_{n+1}=a_n+5\) and \(a_1=4\)

OpenStudy (anonymous):

then \(a_2=a_1+5=4+5=9\) and \[a_3=a_2+5=9+5=14\] and \[a_4=a_3+5=14+5=19\] an in general \[a_{n+1}=a_n+5\]

OpenStudy (anonymous):

Ok think I'm kinda getting iy

OpenStudy (anonymous):

the n plus first term is the next term after the nth term

OpenStudy (anonymous):

Ok but sometimes in these type of problems I see n-1 subscript?

OpenStudy (anonymous):

So the subscript is a way to determine the terms

OpenStudy (anonymous):

makes no difference there is no difference between say \[a_n=a_{n-1}+5\] and \[a_{n+1}=a_n+5\] they both say the next term if 5 more than the previous one

OpenStudy (anonymous):

So it doesn't matter if it is minus or plus

OpenStudy (anonymous):

Oh so it's just moving it to different sides of the equation Is minus always on the sidewith the plus #

OpenStudy (anonymous):

Well Thank you so much u explained this better than my math teacher Ur the best thanks so much

OpenStudy (anonymous):

I'll figure out the rest my own thanks